OpenStudy (anonymous):

Solve. m – 14.6 = 9.74 A. 4.86 B. 23.34 C. 24.34 D. 25.34

OpenStudy (anonymous):

You need to isolate m so to do that add 14.6 to both sides. this cancels the -14.6 on the side of the m and give you m=24.34

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

No problem. glad i could help

OpenStudy (anonymous):

1 more question: Which equation can be used to find the answer? A playground with four sides has a perimeter of 52 ft. Three of the sides have lengths of 9 ft, 16 ft, and 19 ft. What is the length of the fourth side? A. s – 9 + 16 + 19 = 52 s =26 26 ft B. s + 9 + 16 + 19 = 52 s = 8 8 ft C. s – 52 = 9 + 16 – 19 s = 58 58 ft D. s – 9 – 16 – 19 = 52 s = 96 96 ft

OpenStudy (anonymous):

well, when finding perimeter, you add all the sides together, so to find one side you're going to subtract all the given sides from the total perimeter. so: 52-16-19-9=s then simplify. s=8 once you've simplified you can double check by adding 8+19+16+9 together and seeing if it equals 52

OpenStudy (anonymous):

@Xmoses1 plzz help me for this guestion: Which equation can be used to find the answer? A playground with four sides has a perimeter of 52 ft. Three of the sides have lengths of 9 ft, 16 ft, and 19 ft. What is the length of the fourth side? A. s – 9 + 16 + 19 = 52 s =26 26 ft B. s + 9 + 16 + 19 = 52 s = 8 8 ft C. s – 52 = 9 + 16 – 19 s = 58 58 ft D. s – 9 – 16 – 19 = 52 s = 96 96 ft